OpenStudy (anonymous):

An exponential function is: Select one: a. A function that describes a straight line. b. A function of the form x = by, where b is a positive real number other than 1 and x is any real number. c. A function of the form f(x) = bx, where b is a positive real number other than 1 and x is any real number. d. None of the above.

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@mathmath333

OpenStudy (zpupster):

f(x) = x^2 g(x) = 2^x here are a couple of examples of exponential functions.

OpenStudy (anonymous):

ok what else do I do

OpenStudy (anonymous):

ok is it non of the above then

OpenStudy (anonymous):

@zpupster I think the answer is C am I correct

OpenStudy (zpupster):

as written i would say none of the above.

OpenStudy (anonymous):

thnx